OpenStudy (anonymous):

Which expression is equivalent to the given expression? 11(17 – x) A. 11 • 17 – 11x B. 11 • 17 – x C. 17 – 11x D. 11 + 17 – x

OpenStudy (bibby):

When you distribute, you distribute to each term like so: a(b-c) = ab - ac 6(10-4)=(10*6)-(6*4)

OpenStudy (anonymous):

so the answer would be c?

OpenStudy (bibby):

Nope. You need to distribute the 11 to each term. The 11 isn't distributed to the 17 in that one

OpenStudy (anonymous):

A?

OpenStudy (bibby):

yep

OpenStudy (anonymous):

Thanks!